Abigail is 4 times as old as Bonnie. In 6 years, Bonnie will be twice as old as Candice. If, 4 years from now, Abigail will be 36 years old, how old will Candice be in 6 years?
When i was setting an equation for Candice i thought it was 2(B+6) because Bonnie will be twice (2) as old as Candice (b+6), but in the answer key it's (b+6)/2. Can you please explain the whole problem to me ? Thank You

Answer by mananth(16946)

Abigail is 4 times as old as Bonnie. In 6 years, Bonnie will be twice as old as Candice. If, 4 years from now, Abigail will be 36 years old, how old will Candice be in 6 years?
In 4 years Abigail will be 36 years.
So Abigail is 32 years now.
...
let candice age be x now.
after 6 years
she will be x+6 years.
Bonnie will be 2(x+6) years
...
Bonnie's age now will be 2(x+6)-6
4 times Bonnies age = Abigail's age now
4(2(x+6)-6)=32
4(2x+6)=32
8x+24=32
8x=32-24
8x=8
x= 1 Candice's age now
After 6 years years she will be 6+1=7 years